divide 105in to two parts in which the first half part exceeds the one fifth of the second by 35'kindlygive me its step also
Answer by stanbon(75887)

divide 105in to two parts in which the first half part exceeds the one fifth of the second by 35"
-----------------------
x + y = 105"
y-(1/5)x = 35
--------------------
Substitute for "y" and solve for "x":
x + (1/5)x +35 = 105
(6/5)x = 70
x = (5/6)70
---
x = 350/6
x = 58 1/3"
---------------
y = 105 - 58 1/3 = 46 2/3 inches
============================
Cheers,
Stan H.